# How to build transfer matrices, one wave at a time

### v1

@@ 2 General matrix expressions
1D Shear Waves $\dfrac{\omega}{\sqrt{G/\rho}}$ $\sqrt{\mu\rho}$
@@ 3.1 Physical interpretation of propagation in one layer
$T_{11}=\cos\left({\mathrm{k}}\right)={1\over 2}e^{\mathrm{i}{\mathrm{k}x}}+{1\over 2}e^{-\mathrm{i}{\mathrm{k}x}}={1\over 2}e^{\mathrm{i}{\omega x\over c}}+{1\over 2}e^{-\mathrm{i}{\omega x\over c}}\,,$ (5)
@@
So $T_{11}$ would give us the new amplitudes of $f$ created by an impulse, what about $T_{21}$? From the transfer matrix, it follows that it m,ust represent the wave gradient associated to those two new waves in eq. 5:
@@
Once this exercise had revealed the physical meaning of the transfer matrix for one constant piece of medium, nothing deters us from employing the same arguments for more layers.
@@ 3.2 Two-layer transfer matrix
Let us compute $T_{11}$. Introducing $t_{1}=l_{1}/c_{1}$ the time spent to propagate across layer 1 and $t_{2}=l_{2}/c_{2}$ the time spent to propagate acrosss layer 2, we define two possible paths:
@@
    Path I: let us call this path “$[+1,+1]$”, meaning “propagation in 1, transmission in 2”. The phase change $\omega(t_{1}+t_{2})$, proportion of total amplitude $1/\mathsf{T}_{1\to 2}$. It would correspond to
@@
    Path II: let us call this path “$[+1,-1]$”, meaning “propagation in 1, reflection in 2”. Phase change $\omega(t_{1}-t_{2})$, proportion of total amplitude $\mathsf{R}_{1\to 2}/\mathsf{T}_{1\to 2}$. It would correspond to
@@
Expanding sum in the cosines and substituting the coefficients in terms of the impedances in Equation 10, we reach the form that is usually found in the literature:
@@ 3.4 N-layer transfer matrix
Even before presenting here the physical logic that allows building each entry of the transfer matrix path by path, we knew the expressions (23, 25, 24, 26a, 26b) to hold, as they have already been derived in with a purely mathematical tour de force: the first general form of the trace of $\boldsymbol{T}$ was derived exploiting the equivalence between transfer matrices and Möbius transformations [19]. The non-diagonal terms were presented in Ref. [21], and an alternative proof, relying solely on tracking transfer matrix products, has been published in Ref. [20].
@@ 3.5 N-layer scattering matrix
Let us remark that in periodic media, $\mathfrak{Re}(S_{11})=\mathfrak{Re}(S_{22})=\eta$, the half-trace function that governs the dispersion relation in periodic 1D photonic and phoninic crystals. This is possible in periodic systems because one can always shift the unit cell to attain $\mathrm{k}_{1}=\mathrm{k}_{N}$.
@@ 4 Conclusions
We have outlined a combinatorial “path-by-path” procedure to directly construct all the entries of the cumulative transfer matrix $\boldsymbol{T}$ associated with an arbitrary number of layers $N$, and then used these to obtain the corresponding scattering matrix $\boldsymbol{S}$. Their physical meaning was previously concealed behind convoluted mathematical expressions that obscured an intuitive interpretation in terms of the various ways to traverse the layered system and the accompanying changes in amplitude and phase. The path decomposition (formerly termed “harmonic” decomposition) endows transfer matrices with a clear physical significance, so they do not to be regarded just as a convenient computational tool.

### v2

@@ 2 General matrix expressions
1D Shear Waves $\dfrac{\omega}{\sqrt{G/\rho}}$ $\sqrt{G\rho}$
@@ 3.1 Physical interpretation of propagation in one layer
$T_{11}=\cos\left({\mathrm{k}}x\right)={1\over 2}e^{\mathrm{i}{\mathrm{k}x}}+{1\over 2}e^{-\mathrm{i}{\mathrm{k}x}}={1\over 2}e^{\mathrm{i}{\omega x\over c}}+{1\over 2}e^{-\mathrm{i}{\omega x\over c}}\,,$ (5)
@@
So $T_{11}$ would give us the new amplitudes of $f$ created by an impulse, what about $T_{21}$? From the transfer matrix, it follows that it must represent the wave gradient associated to those two new waves in eq. 5:
@@
Once this exercise has revealed the physical meaning of the transfer matrix for one constant piece of medium, nothing deters us from employing the same arguments for more layers.
@@ 3.2 Two-layer transfer matrix
Let us compute $T_{11}$. Introducing $t_{1}=l_{1}/c_{1}$ the time spent to propagate across layer 1 and $t_{2}=l_{2}/c_{2}$ the time spent to propagate across layer 2, we define two possible paths:
@@
    Path I: let us call this path “$[+1,+1]$”, meaning “propagation in 1, transmission in 2”. The phase change would be $\omega(t_{1}+t_{2})$, proportion of total amplitude would be $1/\mathsf{T}_{1\to 2}$. It would correspond to
@@
    Path II: let us call this path “$[+1,-1]$”, meaning “propagation in 1, reflection in 2”. The phase change would be $\omega(t_{1}-t_{2})$, the proportion of total amplitude would be $\mathsf{R}_{1\to 2}/\mathsf{T}_{1\to 2}$. It would correspond to
@@
Expanding the sum in the cosines’ argument and substituting the coefficients in terms of the impedances eq. 7 in Equation 10, we reach the form that is usually found in the literature:
@@ 3.4 N-layer transfer matrix
Even before presenting here the physical logic that allows building each entry of the transfer matrix path by path, we knew the expressions (23, 25, 24, 26a, 26b) to hold, as they have already been derived in a purely mathematical tour de force: the first general form of the trace of $\boldsymbol{T}$ was derived exploiting the equivalence between transfer matrices and Möbius transformations [19]. The non-diagonal terms were presented in Ref. [21], and an alternative proof, relying solely on tracking transfer matrix products, has been published in Ref. [20].
@@ 3.5 N-layer scattering matrix
Let us remark that in periodic media, $\mathfrak{Re}(S_{11})=\mathfrak{Re}(S_{22})=\eta$, the half-trace function that governs the dispersion relation in periodic 1D crystals. This is possible in periodic systems because one can always shift the unit cell to attain $\mathrm{k}_{1}=\mathrm{k}_{N}$.
@@ 4 Conclusions
We have outlined a combinatorial “path-by-path” procedure to directly construct all the entries of the cumulative transfer matrix $\boldsymbol{T}$ associated with an arbitrary number of layers $N$, and then used these to obtain the corresponding scattering matrix $\boldsymbol{S}$. Their physical meaning was previously concealed behind convoluted mathematical expressions that obscured an intuitive interpretation in terms of the various ways to traverse the layered system and the accompanying changes in amplitude and phase. The path decomposition (formerly termed “harmonic” decomposition) endows transfer matrices with a clear physical significance, so they do not have to be regarded just as a convenient computational tool.
Acknowledgements. The support of the Swiss National Science Foundation via Ambizione Grant 216341 “Data-Driven Computational Friction” is gratefully acknowledged.